OpenStudy (anonymous):

-4(2x+3)=16 solve for x & i need the work on how it was done please

OpenStudy (anonymous):

2x+3=16/-4 2x+3=-4 2x=-4-3 2x=-7 x=-7/2 also known as -3 1/2